OpenStudy (anonymous):

At 2:00 p.m. two cars start toward each other from towns 240 miles apart. If the rate of one car is 10 mph faster than the other, how fast does each car go if they meet at 5:00 p.m.? { mph, mph}

OpenStudy (anonymous):

x, x+10 the x's are the speed of the cars & since one is suppose to go 10 miles faster then the other yu add +10. Now you add up their speed, 2x+10 5-2 = 3 hours distance = speed*time 240 = (2x+10)*3 240 = (2x+10)*3 2x+10 = 80 2x = 70 x = 35 x+10 = 45 So the answer is 35 & 45.!